is 0 a solution of 4-3b=4-5b
Answer by Madgamer107(9)

Plug in zero as b to find out.
4-3b = 4-5b
4-3(0) = 4-5(0)
4-0 = 4-0
4 = 4
Since the left and right side do, in fact, equal each other, it can be said that 4 is a possible solution to the equation. It can also be solved algebraically
4 - 3b = 4 - 5b
4 = 4 - 2b
0 = -2b
0 = b